A mass of 1.0 × 103 g of a spring whose force constant is k = 1.0 × 106 dynes / cm is suspended. (to) Determine the period of oscillation, (b) if for t = 0 s the equilibrium shift is 0.5 cm and the velocity + 15 cm / s, determine the x (t)
Answer #1

The period of a spring-mass system is proportional to the square root of the mass and inversely proportional to the square root of the spring constant.

where m is the mass and k is the force constant.

Here, m = 1×103 grams = 1 kilogram

k = 1×106 dynes/cm = 103 N/m

Now, time period,

= 0.199 seconds